Use variation of parameter methods to find the particular solution of xy−(x+1)y+y = x2, given that y1(x) = ex and y2(x) = x + 1 form a fundamental set of solutions for the corresponding homogeneous differential equation.


1
Expert's answer
2022-02-28T11:56:18-0500
"xy''-(x+1)y'+y=x^2"

"y=c_1e^x+c_2(x+1)"

"y'=c_1e^x+c_1'e^x+c_2+c_2'(x+1)"

Let "c_1'e^x+c_2'(x+1)=0." Then


"y'=c_1e^x+c_2"

"y''=c_1e^x+c_1'e^x+c_2'"

Substitute


"c_1xe^x+c_1'xe^x+c_2'x-(x+1)(c_1e^x+c_2)"

"+c_1e^x+c_2(x+1)=x^2"

"c_1'xe^x+c_2'x=x^2"

We have


"c_1'e^x+c_2'(x+1)=0"

"c_1'xe^x+c_2'x=x^2"


"c_1'e^x=-c_2'(x+1)"

"-c_2'x(x+1)+c_2'x=x^2"

"c_2'=-1"

"c_1'=e^{-x}(x+1)"

Then we can take


"c_1=\u2212(x+1)e^{-x}-e^{-x}=\u2212(x+2)e^{-x}"

"c_2=-x"

The particular solution is


"y_p=-x-2-x^2-x"

"y_p=-x^2-2x-2"